# Fornax dwarf spheroidal in MOND: its formation and the survival of its globular clusters

### v1

@@
Michal Bílek 1122    Hongsheng Zhao 22334455
@@
The Fornax dwarf spheroidal galaxy has five massive globular clusters (GCs). They are often used for testing different dark matter and modified gravity theories, because it is difficult to reconcile their old stellar ages with the short time they need to settle in the center of the galaxy due to dynamical friction. Using high resolution $N$-body simulations with the Phantom of Ramses code, we investigate whether the GCs of Fornax can be reconciled with the modified Newtonian dynamics (MOND), namely its QUMOND formulation. Observational data interpreted in MOND indicate that Fornax is a tidal dwarf galaxy formed at redshift $z=0.9$ in a flyby of the Milky Way (MW) and Andromeda galaxies, and that its GCs were initially massive star clusters in the disk of the MW. This helps us to set up and interpret the simulations. In the simulations, a point-mass GC orbits Fornax, and they both orbit the MW. When we ran multiple simulations with varying initial conditions for the GC, we found a 20% probability of Fornax being observed with five unsunk GCs. The unsunk GCs have the observed radial distribution. Moreover, we found: 1) In MOND, Fornax has an orbit around the MW such that the pericenters coincide with the observed peaks in the star formation history of Fornax; 2) The simulations reproduce the observed “diffuse stellar halo” of Fornax; 3) The simulations predict that Fornax has a stellar stream, which could be detectable in the existing data. 4) An extra simulation shows that if Fornax was initially a rotating disky tidal dwarf galaxy, the gravitational influence of the MW would be able to transform it into a nonrotating spheroidal. 5) Sometimes Phantom of Ramses does not conserve angular momentum. This makes the GC sink too fast if it is simulated as an $N$-body object.
@@ 1 Introduction
In the context of Newtonian gravity, various solutions to the mystery of the GCs of Fornax have been proposed. For example, it has been proposed that the dark matter halo of Fornax has a core (Goerdt et al., 2006), the GCs were formed much further from Fornax than currently observed (Angus & Diaferio, 2009), Fornax experienced a major merger (Leung et al., 2020), or that the GCs are embedded in dark matter minihalos (Boldrini et al., 2020). A comprehensive model of galaxy formation in the $\Lambda$CDM cosmology even shows that a few percent of the simulated dwarf spheroidals indeed possess GC systems similar to that of Fornax without any special adjustments (Shao et al., 2021). However, it should be pointed out that regardless of the advances in the problem of the GCs of Fornax, the $\Lambda$CDM cosmology struggles to explain the existence of the disk of satellites of the MW, of which Fornax is clearly a member (see Sect. 2.1 for details).
In this contribution, we investigate the survivability of the GCs of Fornax in modified Newtonian dynamics (MOND, Milgrom, 1983), namely in its QUMOND formulation (Milgrom, 2010). Initial analytic calculations indicated that the problem of survival of the GCs of Fornax (and other dwarf galaxies) is even fiercer in MOND than with Newtonian gravity and dark matter, because of more efficient dynamical friction (Ciotti & Binney, 2004; Sánchez-Salcedo et al., 2006; Nipoti et al., 2008). It has been argued, again from analytic calculations, that the problem can be solved if GCs start their existence near the tidal radius of Fornax (Angus & Diaferio, 2009). Using high-resolution simulations of GCs in ultra-diffuse galaxies, Bílek et al. (2021b) demonstrate that the analytic formulas are correct, but only as long as the GC is not too close to the center of the galaxy. Then the sinking is slowed down by the core-stalling effect (Hernandez & Gilmore, 1998; Read et al., 2006; Petts et al., 2015), which was ignored when deriving the analytical formulas for dynamical friction and is difficult to incorporate into them. That work came to the conclusion that MOND can explain the presence of GCs in ultra-diffuse galaxies. The simulations of Bílek et al. (2024) investigate the survival of GCs in isolated gas-rich galaxies (90% of baryons in gas) both for MOND and $\Lambda$CDM. It turned out that, in MOND, supernova (SN) explosions cause large fluctuations of gravitational potential, which prevent the GCs from sinking, but only if the mass of the GC is less than $4\times 10^{5}\,M_{\odot}$. It remained unclear, however, to what extent this limit is influenced by the numerical implementation of the baryonic processes. There were hints that the rate of SN explosions was smaller than it should.
The main goal of the present work is to investigate the compatibility of QUMOND with Fornax GCs. To that end, we use high-resolution $N$-body simulations of point-mass GCs in a spherical Fornax model without hydrodynamics. Compared to the case of the isolated ultra-diffuse galaxies in Bílek et al. (2024), Fornax has a smaller size, lower mass, and is gravitationally perturbed by the MW. It turns out that, for a proper setup on the simulation and its interpretation, a wider context of the formation of Fornax and its GCs has to be taken into account. That is deduced mostly from observations and partly from older simulations and dynamical models. Apart from answering our main question, the simulations also give many other interesting results, which we also describe here.
Table 1: Parameters of Fornax, its GCs, the MW, and the observing frame.
@@
Heliocentric proper motion $\mu_{\delta}$0.358 mas yr-1 (4)
@@
Stellar masses (0.4, 1.5, 5.0, 0.8, 1.9) ×$\times 10^{5}\,M_{\sun}$ (6)
@@
111(1) Oakes et al. (2022). (2) Intermediate value between McConnachie (2012) (stellar mass obtained from the $V$-band luminosity by assuming the mass-to-light ratio of one) and de Boer et al. (2012) (from a stellar populations synthesis model). (3) Sample S0 of Yang et al. (2022), effective radius converted from the provided Sérsic radius. (4) Battaglia et al. (2022). (5) McConnachie (2012). (6) de Boer & Fraser (2016). (7) Calculated from the coordinates of the GCs by Mackey & Gilmore (2003) and the coordinates and distance of Fornax given above. The distances differ from those given by Mackey & Gilmore (2003) because a different center of Fornax is used. (8) Zhao et al. (2013). (9) GRAVITY Collaboration et al. (2022); Gaia Collaboration et al. (2023). Cartesian coordinates with the center at position of the Sun assumed. The $x$-axis points toward the Galactic center, the $y$-axis in the direction of the motion of the Sun, and the $z$-axis forms a right-handed coordinate system with the other two axes.
@@ 2.5 Fornax was initially gas rich and SN explosions potentially helped the GCs to survive
In this initial work on the subject, we perform only gasless $N$-body simulations. Instead, the influence of star formation is only discussed on the basis of these arguments. We note that Fornax was forming stars vigourously to 2-4 Gyr ago. We assume that up to this point, the influence of SNe was sufficient to push the GCs away from the Fornax center, but only not farther than 0.5-1 kpc, where star formation was happening according to Fig. 3. We assume that from 2 Gyr ago the evolution of the GCs would be the same as in our gasless $N$-body simulations.
@@ 2.6.2 Could the GCs of Fornax have formed in the MW?
Table 3: Statistics of sunk, surviving, and escaped GCs for different assumptions.
@@
 [^2] The “Layer” column gives the radial range within which the GCs start. Ten GCs start from each layer. The “Light (Massive) GC” refers to the $1(5)\times 10^{5}\,M_{\sun}$ GC. The columns $n_{c}$, $n_{s}$, and $n_{e}$ show the numbers of sunk (as for “core”), surviving, and escaped GCs in each layer, respectively. The asterisk symbol indicates that the numbers are assumed, without actually making the simulation. The column $\gamma$ lists the geometric weights of the layers.
@@ 3 Simulations
The main objective of this paper is to investigate whether the observations of the five GCs in Fornax excludes QUMOND, using high-resolution $N$-body gas-less simulations with a spherical Fornax. Unless specified otherwise, each simulation contained the MW, Fornax and one GC.
@@
Figure 4: Orbital decay of the GCs of Fornax, depending on their mass and initial trajectory. The vertical axis denotes the distance between the GG and the center of Fornax. For the light GCs, the two dotted horizontal lines indicate the range of the projected distances of the four lightest GCs of the real Fornax. For the massive GC, the horizontal dotted line indicates the projected distance of the most massive GC of the real Fornax.
@@ 5.3 Diffuse stellar halo of Fornax reproduced
We note that a better agreement cannot be expected because we used simplified initial conditions. The real Fornax at the beginning of the simulated period, was most probably not an $N$-body sphere but a rotating disk with a substantial fraction of gas, see Sects. 2 and 6.
@@ 5.6 MOND-specific deformation of shapes of satellite galaxies
Table 4: Parameters of the disky Fornax in the simulation of tidal stirring.
@@
444(1) Radial density profile is an exponential disk. (2) Vertical density profile is the $\operatorname{sech}^{2}$ profile with a constant scale height. (3) A fixed value of Toomre Q was used across the disk.
@@ 6 Extra simulation: Transformation of a young disky Fornax into a nonrotating spheroid
The simulation of the tidal stirring was the same as the simulations described before (Sect. 3), but Fornax was initially set as a rotating disk without GCs. The initial characteristics of the disky Fornax are listed in Table 4. The disky model was generated using the MOND version (Banik et al., 2020) of the Disk Initial Condition Environment (dice; Perret et al., 2014) code. After some experimentation, we found that tilting the spin of the disk in the direction $(1,0,-1)$ gives good results, which we describe below.
@@ 7 Conclusions
Fornax possesses five classical GCs that are often used for testing dark matter and modified gravity theories. Here we use the GCs to test the QUMOND formulation of MOND using high-resolution $N$-body simulations in the popular PoR code.
@@
  • One should be careful when using PoR. We found that in some settings, the code does not conserve the total angular momentum (Appendix A). For us, when the GC was simulated as a resolved $N$-body object, this caused an overly fast sinking of the GC.
@@ Appendix A Problems with conservation of angular momentum in PoR
Initially, we modeled the GCs as resolved $N$-body objects instead of point masses. In such simulations, they experienced very fast sinking. Eventually, we found that this was because the total angular momentum of all particles in the box is not conserved for the simulations, which indicates a numerical problem. Since this problem has not been described in the literature on PoR yet, we present here more details as a warning.
@@
We developed the following procedure to monitor for the conservation of the angular momentum in a simulation. For each output of the simulation, we first calculated the total angular momentum with respect to the corner of the computational box, ${L}_{\mathrm{raw}}$. This quantity itself is difficult to interpret because the whole system always shows some numerical drift with respect to the computational grid. This numerical drift seems to be an accelerated motion with a lot of random noise. To remove this effect from the evaluation of the conservation of the total internal angular momentum, we also calculated the angular momentum of the center of mass of the system with respect to the corner of the computational box, ${L}_{\mathrm{com}}$. The total internal angular momentum that we are interested in is then ${L}=\@vec{L}_{\mathrm{raw}}-\@vec{L}_{\mathrm{com}}$. We then evaluated the time dependence of the $z$-component of ${L}$, that is $L_{z}$, as a function of time. For an easier interpretation, $L_{z}$ was expressed in the units of the value of $L_{z}$ at the beginning of the simulation, denoted by $L_{z,\mathrm{ini}}$.
@@ Appendix B Influence of the other GCs
To analyze the results, we sorted the different trajectories in these groups: 1) Massive GCs simulated alone; 2) light GCs simulated alone, 3) four light GCs simulated together (i.e., from the reference set), 4) the massive GCs in the simulations with five GCs, 5) the light GCs in the simulations with five GCs. For each group, we calculated for every time the median of distances of the GCs from the center of Fornax within the time window of ±$\pm 250$ Myr. The results are shown in Fig. 13. It shows that the presence of the other GCs does not have a substantial net impact on the sinking of the GCs.
@@
$g_{\mathrm{ext,peri}}\leavevmode\nobreak\ /\leavevmode\nobreak\ a_{0}$ $0.13\pm 0.01$
$g_{\mathrm{ext,apo}}\leavevmode\nobreak\ /\leavevmode\nobreak\ a_{0}$ $0.058\pm 0.001$

### v2

@@
Michal Bílek    Hongsheng Zhao
@@
The Fornax dwarf spheroidal galaxy has five massive globular clusters (GCs). They are often used for testing different dark matter and modified gravity theories, because it is difficult to reconcile their old stellar ages with the short time they need to settle in the center of the galaxy due to dynamical friction. Using high resolution NN-body simulations with the Phantom of Ramses code, we investigate whether the GCs of Fornax can be reconciled with the modified Newtonian dynamics (MOND), namely its QUMOND formulation. Observational data interpreted in MOND indicate that Fornax is a tidal dwarf galaxy formed at redshift $z=0.9$ in a flyby of the Milky Way (MW) and Andromeda galaxies, and that its GCs were initially massive star clusters in the disk of the MW. This helps us to set up and interpret the simulations. In the simulations, a point-mass GC orbits Fornax, and they both orbit the MW. When we ran multiple simulations with varying initial conditions for the GC, we found a 20% probability of Fornax being observed with five unsunk GCs. The unsunk GCs have the observed radial distribution. Moreover, we found: 1) In MOND, Fornax has an orbit around the MW such that the pericenters coincide with the observed peaks in the star formation history of Fornax; 2) The simulations reproduce the observed “diffuse stellar halo” of Fornax; 3) The simulations predict that Fornax has a stellar stream, which could be detectable in the existing data. 4) An extra simulation shows that if Fornax was initially a rotating disky tidal dwarf galaxy, the gravitational influence of the MW would be able to transform it into a nonrotating spheroidal. 5) Sometimes Phantom of Ramses does not conserve angular momentum. This makes the GC sink too fast if it is simulated as an NN-body object.
@@ 1 Introduction
[…]
The main goal of the present work is to investigate the compatibility of QUMOND with Fornax GCs. To that end, we use high-resolution NN-body simulations of point-mass GCs in a spherical Fornax model without hydrodynamics. Compared to the case of the isolated ultra-diffuse galaxies in Bílek et al. (2024), Fornax has a smaller size, lower mass, and is gravitationally perturbed by the MW. It turns out that, for a proper setup on the simulation and its interpretation, a wider context of the formation of Fornax and its GCs has to be taken into account. That is deduced mostly from observations and partly from older simulations and dynamical models. Apart from answering our main question, the simulations also give many other interesting results, which we also describe here.
Table 1: Parameters of Fornax, its GCs, the MW, and the observing frame.111(1) Oakes et al. (2022). (2) Intermediate value between McConnachie (2012) (stellar mass obtained from the $V$-band luminosity by assuming the mass-to-light ratio of one) and de Boer et al. (2012) (from a stellar populations synthesis model). (3) Sample S0 of Yang et al. (2022), effective radius converted from the provided Sérsic radius. (4) Battaglia et al. (2022). (5) McConnachie (2012). (6) de Boer & Fraser (2016). (7) Calculated from the coordinates of the GCs by Mackey & Gilmore (2003) and the coordinates and distance of Fornax given above. The distances differ from those given by Mackey & Gilmore (2003) because a different center of Fornax is used. (8) Zhao et al. (2013). (9) GRAVITY Collaboration et al. (2022); Gaia Collaboration et al. (2023). Cartesian coordinates with the center at position of the Sun assumed. The $x$-axis points toward the Galactic center, the $y$-axis in the direction of the motion of the Sun, and the $z$-axis forms a right-handed coordinate system with the other two axes.
@@
Heliocentric proper motion $\mu_{\delta}$ -0.358 mas yr-1 (4)
@@
Stellar masses (0.4, 1.5, 5.0, 0.8, 1.9) $\times 10^{5}\,M_{\sun}$ (6)
@@ 2.5 Fornax was initially gas rich and SN explosions potentially helped the GCs to survive
In this initial work on the subject, we perform only gasless NN-body simulations. Instead, the influence of star formation is only discussed on the basis of these arguments. We note that Fornax was forming stars vigourously to 2-4 Gyr ago. We assume that up to this point, the influence of SNe was sufficient to push the GCs away from the Fornax center, but only not farther than 0.5-1 kpc, where star formation was happening according to Fig. 3. We assume that from 2 Gyr ago the evolution of the GCs would be the same as in our gasless NN-body simulations.
@@ 2.6.2 Could the GCs of Fornax have formed in the MW?
Table 3: Statistics of sunk, surviving, and escaped GCs for different assumptions.  [^2] The “Layer” column gives the radial range within which the GCs start. Ten GCs start from each layer. The “Light (Massive) GC” refers to the $1(5)\times 10^{5}\,M_{\sun}$ GC. The columns $n_{c}$, $n_{s}$, and $n_{e}$ show the numbers of sunk (as for “core”), surviving, and escaped GCs in each layer, respectively. The asterisk symbol indicates that the numbers are assumed, without actually making the simulation. The column $\gamma$ lists the geometric weights of the layers.
@@ 3 Simulations
The main objective of this paper is to investigate whether the observations of the five GCs in Fornax excludes QUMOND, using high-resolution NN-body gas-less simulations with a spherical Fornax. Unless specified otherwise, each simulation contained the MW, Fornax and one GC.
@@
Figure 4: Orbital decay of the GCs of Fornax, depending on their mass and initial trajectory. The vertical axis denotes the distance between the $G$ and the center of Fornax. For the light GCs, the two dotted horizontal lines indicate the range of the projected distances of the four lightest GCs of the real Fornax. For the massive GC, the horizontal dotted line indicates the projected distance of the most massive GC of the real Fornax.
@@ 5.3 Diffuse stellar halo of Fornax reproduced
We note that a better agreement cannot be expected because we used simplified initial conditions. The real Fornax at the beginning of the simulated period, was most probably not an NN-body sphere but a rotating disk with a substantial fraction of gas, see Sects. 2 and 6.
@@ 5.6 MOND-specific deformation of shapes of satellite galaxies
Table 4: Parameters of the disky Fornax in the simulation of tidal stirring.444(1) Radial density profile is an exponential disk. (2) Vertical density profile is the $\operatorname{sech}^{2}$ profile with a constant scale height. (3) A fixed value of Toomre Q was used across the disk.
@@ 6 Extra simulation: Transformation of a young disky Fornax into a nonrotating spheroid
The simulation of the tidal stirring was the same as the simulations described before (Sect. 3), but Fornax was initially set as a rotating disk without GCs. The initial characteristics of the disky Fornax are listed in Table 4. The disky model was generated using the MOND version (Banik et al., 2020) of the Disk Initial Condition Environment (dice; Perret et al., 2014) code. After some experimentation, we found that tilting the spin of the disk in the direction (1,0,1)(1,0,-1) gives good results, which we describe below.
@@ 7 Conclusions
Fornax possesses five classical GCs that are often used for testing dark matter and modified gravity theories. Here we use the GCs to test the QUMOND formulation of MOND using high-resolution NN-body simulations in the popular PoR code.
@@
  • One should be careful when using PoR. We found that in some settings, the code does not conserve the total angular momentum (Appendix A). For us, when the GC was simulated as a resolved NN-body object, this caused an overly fast sinking of the GC.
@@ Appendix A Problems with conservation of angular momentum in PoR
Initially, we modeled the GCs as resolved NN-body objects instead of point masses. In such simulations, they experienced very fast sinking. Eventually, we found that this was because the total angular momentum of all particles in the box is not conserved for the simulations, which indicates a numerical problem. Since this problem has not been described in the literature on PoR yet, we present here more details as a warning.
@@
We developed the following procedure to monitor for the conservation of the angular momentum in a simulation. For each output of the simulation, we first calculated the total angular momentum with respect to the corner of the computational box, $\@vec{L}_{\mathrm{raw}}$. This quantity itself is difficult to interpret because the whole system always shows some numerical drift with respect to the computational grid. This numerical drift seems to be an accelerated motion with a lot of random noise. To remove this effect from the evaluation of the conservation of the total internal angular momentum, we also calculated the angular momentum of the center of mass of the system with respect to the corner of the computational box, $\@vec{L}_{\mathrm{com}}$. The total internal angular momentum that we are interested in is then $\@vec{L}=\@vec{L}_{\mathrm{raw}}-\@vec{L}_{\mathrm{com}}$. We then evaluated the time dependence of the $z$-component of L\@vec{L}, that is $L_{z}$, as a function of time. For an easier interpretation, $L_{z}$ was expressed in the units of the value of $L_{z}$ at the beginning of the simulation, denoted by $L_{z,\mathrm{ini}}$.
@@ Appendix B Influence of the other GCs
To analyze the results, we sorted the different trajectories in these groups: 1) Massive GCs simulated alone; 2) light GCs simulated alone, 3) four light GCs simulated together (i.e., from the reference set), 4) the massive GCs in the simulations with five GCs, 5) the light GCs in the simulations with five GCs. For each group, we calculated for every time the median of distances of the GCs from the center of Fornax within the time window of $\pm 250$ Myr. The results are shown in Fig. 13. It shows that the presence of the other GCs does not have a substantial net impact on the sinking of the GCs.
@@
$g_{\mathrm{ext,peri}}\penalty 10000\ /\penalty 10000\ a_{0}$ $0.13\pm 0.01$
$g_{\mathrm{ext,apo}}\penalty 10000\ /\penalty 10000\ a_{0}$ $0.058\pm 0.001$